From: China National Space Administration
Posted: Tuesday, May 15, 2001
No change in the status of the POES spacecraft. NOAA-16 rev 03317 / F at 1508Z on May 14: A monitor-only pass during which HRPT was supposed to be recorded by the IMUX for post-pass transfer was unsuccessful. Numerous attempts were made to accomplish the post-pass playback from the IMUX, but the CDA concluded that the IMUX did not record the data. NOAA-11 SBUV flex memory three will be re-loaded to include TOMS optimized and total ozone wavelengths. This will take place during the 1504z pass on Friday, May 18. POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft. Returned OPS49 to the nominal SCP 1 PRADS / SCP 2 Monitor attitude mode, loaded the SCRUB flight software package ID and enabled the SCRUB software on rev 7337.
OPS47's flight software issued a basic mode request due to the number of lost star transits reaching the threshold limit. Initial analysis indicate the high number of lost star transits was due to environmental conditions. DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft. GOES operations were nominal over the past 24 hours.
Questions or comments to Tim.Walsh@noaa.gov
// end //